# HPC-Driven Modeling with ML-Based Surrogates for Magnon-Photon Dynamics in Hybrid Quantum Systems

### v1

@@ 2 Results
We evaluate the performance of our proposed ML model for modeling long-horizon dynamical fields under two scenarios: predictions at a single spatial location (Figure 4); and multiple probing locations (Figure 5). The model is trained using limited input data and physics constraint represented by a specific physical loss term (refer to Physics Loss. in Methods for details). A representative example is shown in Figure 4(a), where the ML model is given only 200 temporal samples (0.6 ns) as input and predicts the long subsequent 1480 temporal samples (4.1 ns) under a DC magnetic bias of 2050 Oersted. The frequency spectrum in Figure 4(b) and Figure 4(d) shows that the ML surrogate reproduces the hybridized polaritons’ characteristics, confirming it has learned the underlying physics rather than simply fitting short‑term temporal trends. An ablation test (Fig. 4(c)) shows physics‑informed training achieves lower long‑horizon error than purely data‑driven models, while requiring fewer samples and training steps. Using the ESPRIT algorithm [61], we extracted the two polariton frequencies and their QQ‑factors directly from the time traces in Figure 4(a); the results are plotted in panels (e) and (f). Frequencies match almost exactly, while small QQ‑factor offsets persist—typical for high‑$Q$ systems where tiny signal errors and hyper‑parameter choices strongly affect the fitted decay rate. Even so, the physics‑informed model reproduces both resonant modes with high fidelity, confirming its accuracy for precision‑sensitive metrics.
@@
Figure 4: Physics-informed ML time sequence prediction at one single spatial probing point. (a) Time-series of the probed $M_{x}$ field: ML prediction versus numerical ground truth. (b) Fourier spectra of the same traces as (a). (c) The root mean squared error (RMSE) as a function of forcast length, with and without the physical loss. (d) Magnitude spectra of $M_{x}$ under varying magnetic DC bias $H_{0}$. The x-axis is the bias magnetic field in Oersted, and the y-axis is the frequency in GHz; color indicates field amplitude. (e) Hybrid-mode eigen frequecies extracted with ESPRIT [61]-calculated eigen frequencies of the hybrid modes. (f) Corresponding ESPRIT-calculated $Q$ factors of the hybrid modes.
@@ 3 Discussion
We demonstrated hybrid quantum simulation using a magnon-photon testbed, yet the framework is readily extensible to other hybrid magnonic, cavity electrodynamic, and phononic systems  [9, 62, 63, 64]. Ongoing work therefore targets a generalized engine for quantum communication, storage, and computation that natively supports these additional interactions. Our current ML surrogate treats each mesh point as an independent one-dimensional time series, limiting its view of the device’s collective field behavior. We therefore aim to develop spatiotemporal models that learn the coupled evolution of two- and three-dimensional fields, preserve spatial coherence, and predict the full device response with high fidelity.
@@ 4 Methods
where $\bf{E}$} is the electric field, $\bf{B}$} the magnetic flux density, $\bf{H}$} the magnetic-field intensity, and ${\bf{J}}_{\rm src}$ an impressed current source. Material parameters are the conductivity $\sigma$, vacuum permittivity $\epsilon_{0}$, relative permittivity $\epsilon_{r}$, and vacuum permeability $\mu_{0}$. The constitutive law linking $\bf{H}$} and $\bf{B}$} is ${\bf{B}}=\mu{\bf{H}}=\mu_{0}({\bf{H}}+{\bf{M}}),$ with $\bf{M}$} the magnetization density and $\mu=\mu_{0}\mu_{r}$ the permeability. This makes Equation 4 into $\nabla\times{\bf{E}}=-{\partial{\bf{H}}}/{\partial t}-{\partial{\bf{M}}}/{\partial t}$. For non-magnetic media, $\mu_{r}\!\approx\!1$ since ${\bf{M}}\!\approx\!0$. In ferromagnets, $\bf{M}$} must be evolved (e.g., via the Landau-Lifshitz-Gilbert (LLG) equation) to capture magnon dynamics. The LLG equation is expressed as:
@@
where $\gamma=-1.759\times 10^{11}$ C/kg is the gyromagnetic ratio and $\alpha$ is the dimensionless Gilbert damping constant ($10^{-5}$ to $10^{-3}$ in typical ferromagnets). The effective field $\bf{H}_{\rm eff}$} collects all magnetic interactions. Here we retain two terms essential for magnon-photon coupling: the time-dependent EM field ${\bf{H}}_{\mathrm{EM}}$, and the static bias ${\bf{H}}_{0}$. Because the ferromagnet is driven close to saturation, the magnetization magnitude is taken constant, $|{\bf{M}}|=M_{s}$, where $M_{s}$ is the (spatially dependent) saturation magnetization. We employ an iterative scheme that uses a trapezoidal discretization of the vector form of the LLG equation (5) described in [29], to solve for ${\bf{M}}^{n+1,r}$, where $n$ denotes the time step and $r$ is the counter index for the iterative procedure:
@@
We build our solver on AMReX [56], the DOE Exascale Computing Project’s block-structured adaptive mesh refinement (AMR) library. AMReX provides structured-grid PDE infrastructure with performance portability: the domain is decomposed into non-overlapping boxes (“grids”) that are distributed across MPI ranks, and execution follows an MPI+X model: X=OpenMP on multicore CPUs; or X=CUDA on NVIDIA GPUs. A single front-end code path uses AMReX loop abstractions, which select the appropriate back end at compile time. Each rank updates only the data it owns using triply nested cell loops. In the pure-MPI case, these are standard i-j-k loops; with MPI+OpenMP the loops are further tiled and assigned to threads; with MPI+CUDA AMReX launches device kernels that map cells to CUDA threads. AMReX manages data residency so arrays remain on the GPU whenever possible, minimizing host-device traffic; inter-rank communication and host/GPU transfers are therefore limited to ghost-cell exchanges performed a small number of times per time step. The efficiency and scalability of our code using MPI+CUDA simulations on NERSC Perlmutter systems is shown in Figure 6. Data from the simulation can be efficiently written using a user-defined number of MPI ranks, to prevent overwhelming the system with simultaneous writes. Visualization was performed with the publicly-available software package, VisIt [65].
@@
We use the Long Expressive Memory (LEM) model [57] as a robust approach for modeling complex temporal dynamics. LEM is specifically designed to capture multi-scale and long-term behaviors in systems characterized by intricate temporal structures. Its architecture offers a mathematical foundation for handling sequences with varying temporal scales, while ensuring gradient stability and high expressivity. EM fields, which exhibit a combination of rapid oscillations (e.g., high-frequency wave propagation) and slower dynamics (e.g., energy dissipation), present a challenging time-series modeling task. The LEM framework is well-suited for our scientific task due to the capability to effectively learn and predict diverse and interdependent behaviors within EM dynamics. The LEM model is rooted in multi-scale Ordinary Differential Equations (ODEs) that can be used to capture fast and slow dynamics. Let us consider a general formulation of a two-scale ODE system [57], given as
@@
Physics-informed ML has been shown to effectively preserve physical properties in various domains [39]. Physics-informed loss function is adopted into the LEM framework to further improve the model’s ability to learn the underlying physics. We adopted the small-signal approximation (the dynamic magnetization in the bias direction is negligible, i.e. $M_{y}\approx 0$) in the LLG equation when deriving the physics loss term, which is
@@
The analytical derivatives $M_{x}(\textrm{LLG})$ and $M_{z}(\textrm{LLG})$ are the $x$- and $z$-components derived from the trapezoidal discretization of the LLG equation, as given in Eqs. (6)-(8). Given a sequence of network predictions for $n=0,\dots,T$, these residuals are evaluated for $n=0,\dots,T{-}1$.
From the ML side, we estimate the time derivatives from the predicted sequence using a finite forward-Euler discretization:
@@ 7 Acknowledgments
This work was supported by Laboratory Directed Research and Development (LDRD) funding from Berkeley Lab, provided by the Director, Office of Science, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This work was also supported by the US Department of Energy, Office of Science, the Advanced Scientific Computing Research (ASCR) program, the Quantum Systems Accelerator (QSA) Research Center, under contract no. DE-AC02-05-CH11231. This work was also supported by the US Department of Energy, Office of Science, the Microelectronics Science Research Center, under contract no. DE-AC02-05-CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231. The work at the University of Wisconsin-Madison was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award Number DE-SC0020145 as part of the Computational Materials Sciences Program (Y.Z. and J.-M.H.).
@@ 1 Analytical model of a 1D hybrid magnon–photon cavity
Consider a 1D hybrid magnon-photon cavity where all physical quantities vary only along the $z$ axis; see Fig. X.

### v2

@@ 2 Results
We evaluate the performance of our proposed ML model for modeling long-horizon dynamical fields under two scenarios: predictions at a single spatial location (Figure 4); and multiple probing locations (Figure 5). The model is trained using limited input data and physics constraint represented by a specific physical loss term (refer to Physics Loss. in Methods for details). A representative example is shown in Figure 4(a), where the ML model is given only 200 temporal samples (0.6 ns) as input and predicts the long subsequent 1480 temporal samples (4.1 ns) under a DC magnetic bias of 2050 Oersted. The frequency spectrum in Figure 4(b) and Figure 4(d) shows that the ML surrogate reproduces the hybridized polaritons’ characteristics, confirming it has learned the underlying physics rather than simply fitting short‑term temporal trends. An ablation test (Fig. 4(c)) shows physics‑informed training achieves lower long‑horizon error than purely data‑driven models, while requiring fewer samples and training steps. Using the ESPRIT algorithm [esprit], we extracted the two polariton frequencies and their QQ‑factors directly from the time traces in Figure 4(a); the results are plotted in panels (e) and (f). Frequencies match almost exactly, while small QQ‑factor offsets persist—typical for high‑$Q$ systems where tiny signal errors and hyper‑parameter choices strongly affect the fitted decay rate. Even so, the physics‑informed model reproduces both resonant modes with high fidelity, confirming its accuracy for precision‑sensitive metrics.
@@
Figure 4: Physics-informed ML time sequence prediction at one single spatial probing point. (a) Time-series of the probed $M_{x}$ field: ML prediction versus numerical ground truth. (b) Fourier spectra of the same traces as (a). (c) The root mean squared error (RMSE) as a function of forcast length, with and without the physical loss. (d) Magnitude spectra of $M_{x}$ under varying magnetic DC bias $H_{0}$. The x-axis is the bias magnetic field in Oersted, and the y-axis is the frequency in GHz; color indicates field amplitude. (e) Hybrid-mode eigen frequecies extracted with ESPRIT [esprit]-calculated eigen frequencies of the hybrid modes. (f) Corresponding ESPRIT-calculated $Q$ factors of the hybrid modes.
@@ 3 Discussion
We demonstrated hybrid quantum simulation using a magnon–photon testbed, yet the framework is readily extensible to other hybrid magnonic, cavity electrodynamic, and phononic systems  [Chen2018, Bossini2021, Hwang2024, Aspelmeyer2014]. Ongoing work therefore targets a generalized engine for quantum communication, storage, and computation that natively supports these additional interactions. Our current ML surrogate treats each mesh point as an independent one-dimensional time series, limiting its view of the device’s collective field behavior. We therefore aim to develop spatiotemporal models that learn the coupled evolution of two- and three-dimensional fields, preserve spatial coherence, and predict the full device response with high fidelity.
@@ 4 Methods
where $\bf{E}$} is the electric field, $\bf{B}$} the magnetic flux density, $\bf{H}$} the magnetic-field intensity, and ${\bf{J}}_{\rm src}$ an impressed current source. Material parameters are the conductivity $\sigma$, vacuum permittivity $\epsilon_{0}$, relative permittivity $\epsilon_{r}$, and vacuum permeability $\mu_{0}$. The constitutive law linking $\bf{H}$} and $\bf{B}$} is ${\bf{B}}=\mu{\bf{H}}=\mu_{0}({\bf{H}}+{\bf{M}}),$ with $\bf{M}$} the magnetization density and $\mu=\mu_{0}\mu_{r}$ the permeability. This makes Equation 4 into $\nabla\times{\bf{E}}=-{\partial{\bf{H}}}/{\partial t}-{\partial{\bf{M}}}/{\partial t}$. For non-magnetic media, $\mu_{r}\!\approx\!1$ since ${\bf{M}}\!\approx\!0$. In ferromagnets, $\bf{M}$} must be evolved (e.g., via the Landau–Lifshitz–Gilbert (LLG) equation) to capture magnon dynamics. The LLG equation is expressed as:
@@
where $\gamma=-1.759\times 10^{11}$ C/kg is the gyromagnetic ratio and $\alpha$ is the dimensionless Gilbert damping constant ($10^{-5}$ to $10^{-3}$ in typical ferromagnets). The effective field $\bf{H}_{\rm eff}$} collects all magnetic interactions. Here we retain two terms essential for magnon-photon coupling: the time-dependent EM field ${\bf{H}}_{\mathrm{EM}}$, and the static bias ${\bf{H}}_{0}$. Because the ferromagnet is driven close to saturation, the magnetization magnitude is taken constant, $|{\bf{M}}|=M_{s}$, where $M_{s}$ is the (spatially dependent) saturation magnetization. We employ an iterative scheme that uses a trapezoidal discretization of the vector form of the LLG equation (5) described in [Yao2022], to solve for ${\bf{M}}^{n+1,r}$, where $n$ denotes the time step and $r$ is the counter index for the iterative procedure:
@@
We build our solver on AMReX [AMReX_JOSS], the DOE Exascale Computing Project’s block-structured adaptive mesh refinement (AMR) library. AMReX provides structured-grid PDE infrastructure with performance portability: the domain is decomposed into non-overlapping boxes (“grids”) that are distributed across MPI ranks, and execution follows an MPI+X model: X=OpenMP on multicore CPUs; or X=CUDA on NVIDIA GPUs. A single front-end code path uses AMReX loop abstractions, which select the appropriate back end at compile time. Each rank updates only the data it owns using triply nested cell loops. In the pure-MPI case, these are standard i-j-k loops; with MPI+OpenMP the loops are further tiled and assigned to threads; with MPI+CUDA AMReX launches device kernels that map cells to CUDA threads. AMReX manages data residency so arrays remain on the GPU whenever possible, minimizing host–device traffic; inter-rank communication and host/GPU transfers are therefore limited to ghost-cell exchanges performed a small number of times per time step. The efficiency and scalability of our code using MPI+CUDA simulations on NERSC Perlmutter systems is shown in Figure 6. Data from the simulation can be efficiently written using a user-defined number of MPI ranks, to prevent overwhelming the system with simultaneous writes. Visualization was performed with the publicly-available software package, VisIt [VisIt].
@@
We use the Long Expressive Memory (LEM) model [rusch2022iclr] as a robust approach for modeling complex temporal dynamics. LEM is specifically designed to capture multi-scale and long-term behaviors in systems characterized by intricate temporal structures. Its architecture offers a mathematical foundation for handling sequences with varying temporal scales, while ensuring gradient stability and high expressivity. EM fields, which exhibit a combination of rapid oscillations (e.g., high-frequency wave propagation) and slower dynamics (e.g., energy dissipation), present a challenging time-series modeling task. The LEM framework is well-suited for our scientific task due to the capability to effectively learn and predict diverse and interdependent behaviors within EM dynamics. The LEM model is rooted in multi-scale Ordinary Differential Equations (ODEs) that can be used to capture fast and slow dynamics. Let us consider a general formulation of a two-scale ODE system [rusch2022iclr], given as
@@
Physics-informed ML has been shown to effectively preserve physical properties in various domains [karniadakis2021physics]. Physics-informed loss function is adopted into the LEM framework to further improve the model’s ability to learn the underlying physics. We adopted the small-signal approximation (the dynamic magnetization in the bias direction is negligible, i.e. $M_{y}\approx 0$) in the LLG equation when deriving the physics loss term, which is
@@
The analytical derivatives $M_{x}(\textrm{LLG})$ and $M_{z}(\textrm{LLG})$ are the $x$- and $z$-components derived from the trapezoidal discretization of the LLG equation, as given in Eqs. (6)–(8). Given a sequence of network predictions for $n=0,\dots,T$, these residuals are evaluated for $n=0,\dots,T{-}1$.
From the ML side, we estimate the time derivatives from the predicted sequence using a finite forward–Euler discretization:
@@ 7 Acknowledgments
This work was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research Program, the Quantum Systems Accelerator Research Center, and the Microelectronics Science Research Center, under contract No. DE-AC02-05-CH11231. This work was also supported in part by previous breakthroughs obtained through the Laboratory Directed Research and Development Program of Lawrence Berkeley National Laboratory under U.S. Department of Energy Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231. The work at the University of Wisconsin-Madison was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award Number DE-SC0020145 as part of the Computational Materials Sciences Program (Y.Z. and J.-M.H.).
@@ 1 Analytical model of a 1D hybrid magnon–photon cavity
Consider a 1D hybrid magnon-photon cavity where all physical quantities vary only along the $z$ axis; see Supplementary Fig. 1.